OpenStudy (anonymous):

-5w^8 * 9w need help and I will medal .

OpenStudy (anonymous):

@desireesmith15

Directrix (directrix):

-5w^8 * 9w ----------- What is negative 5 times 9 =

OpenStudy (anonymous):

-45

Directrix (directrix):

-5w^8 * 9w = -5 * 9 * w^8 * w^1

Directrix (directrix):

What does this equal? w^8 * w^1 =

OpenStudy (anonymous):

9

Directrix (directrix):

That is the exponent. What raised to the 9th power?

OpenStudy (anonymous):

-45w^9

Directrix (directrix):

Correct.

OpenStudy (anonymous):

yaya

OpenStudy (anonymous):

help with this one to W^2 *(-7w^9)

Directrix (directrix):

See what you can do with it, okay.